[figure: 9.7 cm]

THERMODYNAMICS OF PSEUDOSPIN-ELECTRON MODEL

AT WEAK COUPLING

I. V. Stasyuk, T. S. Mysakovych Institute for Condensed Matter Physics of the National Academy of Science of Ukraine, 1 Svientsitskii St., UA-79011 Lviv, Ukraine

The thermodynamically equilibrium states and phase transitions in the pseudospinelectron model (PEM) with transverse field are studied in the case of weak coupling. The model appeared in connection with the investigation of the role of the anharmonic structure elements in properties of the high-T_{c} superconductors. On the other hand, the PEM is in a close relation to the Falicov-Kimball (FK) model for which the variety of homogeneous, modulated and incommensurate phases as well as the possibility of phase separation are known. In comparison with PEM there exists a difference in the regime of thermodynamical averaging. Besides, the term with transverse field which is responsible for the proper dynamics in the localized particle subsystem is absent in the FK model; thus, the thermodynamics of the PEM can have its own features. In the μ = const regime the first order phase transitions within the uniform phase as well as the first or the second order transitions into the chess-board or incommensurate phase are revealed depending on the location of the chemical potential. In the regime n = const it corresponds to the possibility of separation into phases of the different type. We have also investigated the role of the tunneling-like splitting: besides the general decrease of the phase transition temperatures and narrowing of the segregation regions, it influences the conditions of appearance of the modulated or incommensurate phases.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : I. V. Stasyuk

Address for correspondence : Institute for Condensed Matter Physics of the National Academy of Science of Ukraine, 1 Svientsitskii St., UA-79011 Lviv, Ukraine

Email address : ista@icmp.lviv.ua